# ReAcTree: Hierarchical LLM Agent Trees with Control Flow for Long-Horizon Task Planning

### v1

@@ 5.1. Experimental Setup
Datasets and Simulators. We follow the evaluation protocol of LoTa-Bench (Choi et al., 2024), which provides two dataset-simulator pairs for language-based task planning: WAH-NL (Puig et al., 2021; Choi et al., 2024) with VirtualHome (Puig et al., 2018), and ALFRED (Shridhar et al., 2020) with AI2THOR (Kolve et al., 2017). We extend both environments to more realistic, partially observable settings, where the agent receives limited textual observations after each action. All objects additionally include class and instance identifiers for precise grounding. We conduct primary experiments on WAH-NL with VirtualHome, as it features longer-horizon, multi-room tasks with multiple subgoals, and use ALFRED with AI2THOR, which contains relatively short-horizon, single-room, single-goal tasks, mainly for complementary validation.
WAH-NL, an extension of WAH (Puig et al., 2021), comprises 250 training and 100 test tasks across five categories. We use the training set to build episodic memory and the test set for evaluation. To improve reliability, we manually corrected four test tasks with mismatched instructions and goal conditions (see Appendix B.1 for details). ALFRED includes seven task types. In LoTa-Bench, the pick and place two objects task was omitted due to the absence of instance identifiers. We include this task in our evaluation by incorporating simulator-provided instance identifiers. As with WAH-NL, the training set is used to build episodic memory, and evaluation is conducted on the valid-seen and valid-unseen splits.
@@
To bootstrap episodic memory, we first manually collected a small number of trajectories and used them as in-context examples to run ReAcTree with LLaMA-3.1 70B (Dubey et al., 2024) over the training set. Only successful executions were retained in episodic memory. For WAH-NL, we collected one manual trajectory per task type and executed the full training set. For ALFRED, we collected three per type and stored up to 100 successful trajectories per type to reduce computational overhead, given the large training set (21K tasks).
The retrieval size was constrained such that the total length of retrieved trajectories used as in-context examples did not exceed 5K tokens. The decision cap was set to 200 for WAH-NL and 100 for ALFRED. We used the Guidance library (Microsoft, 2023) for text generation, employing temperature 0.0 for free-form generation and leveraging its constrained generation capabilities to deterministically select acting actions and control flow types. Prompt templates for both WAH-NL and ALFRED are available in Appendix E.1.
Baselines. We compare ReAcTree and ReAcTree+WM against five baselines: ZSP (Huang et al., 2022a), Tree-Planner (Hu et al., 2024) with $N=25$ or $50$ sampled plans, ReAct (Yao et al., 2023), and its working memory variant ReAct+WM. All baselines are evaluated on WAH-NL using the same retrieval size and decision cap as ReAcTree. For ALFRED, we report results only for ReAct+WM, as it serves as the strongest baseline on WAH-NL. Further implementation details and prompt templates are provided in Appendices C and E.2, respectively.
@@ 5.2. Main Results
We first present the primary evaluation results on the WAH-NL dataset, which features complex, long-horizon tasks under partial observability. Table 1 compares five baselines (ZSP, Tree-Planner with $N=25$ or $50$, ReAct, ReAct+WM) and our methods (ReAcTree, ReAcTree+WM) across seven LLMs ranging from small (7–14B) to large (70–72B), including LLaMA 3.1 (Dubey et al., 2024), Qwen 2.5 (Yang et al., 2024), Mistral (Jiang et al., 2023), Gemma 2 (Team et al., 2024), and Phi-4-reasoning-plus (Abdin et al., 2025) (see the Appendix D for the full list of language models and their HuggingFace identifiers). Both GSR and SSR are reported.
@@
To demonstrate how ReAcTree plans and executes complex tasks, we analyze a representative task: Make sure there is a wine and a juice on the coffee table. As shown in Figure 3, ReAcTree+WM successfully solves it by decomposing into two subgoals: move the wine onto the coffee table and move the juice onto the coffee table, and executing them in parallel. It explores multiple rooms using fallback strategies, whereas ReAct+WM remains confined to kitchen 1, failing to locate the wine in bedroom 1. Full trajectories are provided in Appendix G, and the corresponding failure case for ReAct+WM is visualized in Appendix F.
@@
Beyond quantitative gains, ReAcTree+WM also handles complex tasks that require more precise procedures. For instance, given the instruction place a cooked potato slice in the fridge, both methods slice the potato, but ReAct+WM skips the cooking step and places it directly in the fridge, whereas ReAcTree+WM correctly heats it using the microwave before storage by decomposing the instruction into four subgoals with a sequence control flow node: find and pick up knife, slice and pick up potato, cook and pick up potato, and place potato in fridge. The execution trajectories for both methods are visualized in Appendix H.
@@ 5.3. Memory Ablation
All subsequent experiments in this section are conducted on the WAH-NL dataset, featuring long-horizon, multi-room tasks. We study the impact of episodic memory (EM) and working memory (WM) through ablation experiments with Qwen 2.5 models (7B and 72B). When EM is disabled, we use the manually annotated trajectory of a randomly selected training task as a fixed in-context example for all test tasks. For ReAcTree, trajectories from all agent nodes within the selected task are concatenated into a single in-context example. Table 3 reports GSR and SSR of ReAct and ReAcTree under four memory configurations: no memory, WM only, EM only, and EM+WM.
@@
In contrast, the 72B model is robust to this degradation. Its performance improves significantly with WM alone (+16%p GSR for ReAcTree), as its greater reasoning capacity allows it to handle the additional complexity gracefully, even with suboptimal in-context examples.
@@ 5.5. Computational Cost Analysis
To analyze the computational cost of ReAcTree, we conducted a controlled comparison. We measured three representative configurations using LLaMA 3.1: ReAct+WM (70B), ReAcTree+WM (70B), and ReAcTree+WM (8B). All methods were evaluated on the same 19 tasks that all three configurations successfully completed. For fairness in execution time comparison, all experiments were run on identical hardware with two H100 GPUs and one H200 GPU.
Our analysis considers two perspectives: (1) the trade-off between performance and efficiency, and (2) token-based resource usage as a proxy for GPU memory demand.
@@
Table 5. Average execution time, decision steps, and goal/subgoal success rates (GSR/SSR) for three configurations on shared successful tasks.
@@
Table 6. Peak and average token usage per decision across three configurations.
@@ 6. Conclusion
We introduced ReAcTree, a hierarchical task planner that extends ReAct by dynamically constructing an agent tree in the subgoal space. Each agent node can reason, act, or expand the tree with new subgoals, while control flow nodes coordinate their execution through sequence, fallback, or parallel strategies. This design prevents the error accumulation typical of single-pass trajectories and makes long-horizon tasks more tractable. To strengthen in-context learning and coordination, we introduced episodic memory, which stores subgoal-level experiences for retrieval, and working memory, which shares environment-specific observations across nodes. Experiments on WAH-NL and ALFRED under partial observability show that ReAcTree consistently outperforms strong baselines across diverse LLMs, demonstrating improved reliability, generalization to unseen environments, and even enabling smaller models to rival larger ones.

### v2

@@ 5.1. Experimental Setup
Datasets and Simulators. We follow the evaluation protocol of LoTa-Bench (choi2024lotabench), which provides two dataset-simulator pairs for language-based task planning: WAH-NL (puig2021watchandhelp; choi2024lotabench) with VirtualHome (puig2018virtualhome), and ALFRED (shridhar2020alfred) with AI2THOR (kolve2017ai2). We extend both environments to more realistic, partially observable settings, where the agent receives limited textual observations after each action. All objects additionally include class and instance identifiers for precise grounding. We conduct primary experiments on WAH-NL with VirtualHome, as it features longer-horizon, multi-room tasks with multiple subgoals, and use ALFRED with AI2THOR, which contains relatively short-horizon, single-room, single-goal tasks, mainly for complementary validation.
WAH-NL, an extension of WAH (puig2021watchandhelp), comprises 250 training and 100 test tasks across five categories. We use the training set to build episodic memory and the test set for evaluation. To improve reliability, we manually corrected four test tasks with mismatched instructions and goal conditions (see Appendix B.1 for details). ALFRED includes seven task types. In LoTa-Bench, the pick and place two objects task was omitted due to the absence of instance identifiers. We include this task in our evaluation by incorporating simulator-provided instance identifiers. As with WAH-NL, the training set is used to build episodic memory, and evaluation is conducted on the valid-seen and valid-unseen splits.
@@
To bootstrap episodic memory, we first manually collected a small number of trajectories and used them as in-context examples to run ReAcTree with LLaMA-3.1 70B (dubey2024llama) over the training set. Only successful executions were retained in episodic memory. For WAH-NL, we collected one manual trajectory per task type and executed the full training set. For ALFRED, we collected three per type and stored up to 100 successful trajectories per type to reduce computational overhead, given the large training set (21K tasks).
The retrieval size was constrained such that the total length of retrieved trajectories used as in-context examples did not exceed 5K tokens. The decision cap was set to 200 for WAH-NL and 100 for ALFRED. We used the Guidance library (guidance) for text generation, employing temperature 0.0 for free-form generation and leveraging its constrained generation capabilities to deterministically select acting actions and control flow types. Prompt templates for both WAH-NL and ALFRED are available in Appendix E.1.
Baselines. We compare ReAcTree and ReAcTree+WM against five baselines: ZSP (huang2022language), Tree-Planner (hu2024treeplanner) with $N=25$ or $50$ sampled plans, ReAct (yao2023react), and its working memory variant ReAct+WM. All baselines are evaluated on WAH-NL using the same retrieval size and decision cap as ReAcTree. For ALFRED, we report results only for ReAct+WM, as it serves as the strongest baseline on WAH-NL. Appendices C and E provide implementation details and prompts.
@@
[…]
A success case of ReAcTree in the WAH-NL simulator. The figure displays step-by-step agent actions for a complex task, accompanied by the generated behavior tree structure and a table listing the specific subgoal for each node.
@@ 5.2. Main Results
We first present the primary evaluation results on the WAH-NL dataset, which features complex, long-horizon tasks under partial observability. Table 1 compares five baselines (ZSP, Tree-Planner with $N=25$ or $50$, ReAct, ReAct+WM) and our methods (ReAcTree, ReAcTree+WM) across seven LLMs ranging from small (7–14B) to large (70–72B), including LLaMA 3.1 (dubey2024llama), Qwen 2.5 (yang2024qwen2), Mistral (jiang2023mistral), Gemma 2 (team2024gemma), and Phi-4-reasoning-plus (abdin2025phi) (see Appendix D for the full list). Both GSR and SSR are reported.
@@
To demonstrate how ReAcTree plans and executes complex tasks, we analyze a representative task: Make sure there is a wine and a juice on the coffee table. As shown in Figure 3, ReAcTree+WM successfully solves it by decomposing into two subgoals: move the wine onto the coffee table and move the juice onto the coffee table, and executing them in parallel. It explores multiple rooms using fallback strategies, whereas ReAct+WM remains confined to kitchen 1, failing to locate the wine in bedroom 1. Full trajectories are in Appendix G, and the corresponding ReAct+WM failure case is visualized in Appendix F.
@@
Beyond quantitative gains, ReAcTree+WM also handles complex tasks requiring more precise procedures. For instance, given the instruction place a cooked potato slice in the fridge, both methods slice the potato, but ReAct+WM skips the cooking step and places it directly in the fridge, whereas ReAcTree+WM correctly heats it using the microwave before storage by decomposing the instruction into four subgoals with a sequence control flow node: find and pick up knife, slice and pick up potato, cook and pick up potato, and place potato in fridge (see Appendix H for trajectories of both methods).
@@ 5.3. Memory Ablation
All subsequent experiments in this section are conducted on the WAH-NL dataset, featuring long-horizon, multi-room tasks. We study the impact of episodic memory (EM) and working memory (WM) through ablations with Qwen 2.5 models (7B and 72B). When EM is disabled, we use the manually annotated trajectory of a randomly selected training task as a fixed in-context example for all test tasks. For ReAcTree, trajectories from all agent nodes within the selected task are concatenated into a single in-context example. Table 3 reports GSR and SSR of ReAct and ReAcTree under four configurations: no memory, WM only, EM only, and EM+WM.
@@
In contrast, the 72B model is robust to this degradation. Its performance improves significantly with WM alone (+16%p GSR for ReAcTree), as its greater reasoning capacity handles the additional complexity gracefully, even with suboptimal in-context examples.
@@ 5.5. Computational Cost Analysis
To analyze computational cost of ReAcTree, we compared three representative configurations using LLaMA 3.1: ReAct+WM (70B), and ReAcTree+WM (70B, 8B). All methods were evaluated on the same 19 tasks successfully completed by all configurations. For fairness, all experiments were run on identical hardware with two H100 GPUs and one H200 GPU. Our analysis considers two perspectives: (1) the trade-off between performance and efficiency, and (2) token-based resource usage as a proxy for GPU memory demand.
@@
Table 5. Average execution time, decision steps, and GSR/SSR on shared successful tasks.
@@
Table 6. Peak and average token usage per decision.
@@ 5.6. Failure Case Analysis
[…]
A Sankey diagram showing the breakdown of failure cases into four main categories (Search, Execution, Ambiguous, Expand) and their respective sub-types.
@@ 6. Conclusion
We introduced ReAcTree, a hierarchical task planner that extends ReAct by dynamically constructing an agent tree in subgoal space. Each agent node can reason, act, or expand the tree with new subgoals, while control flow nodes coordinate their execution through sequence, fallback, or parallel strategies. This design prevents error accumulation typical of single-pass trajectories and makes long-horizon tasks more tractable. To strengthen in-context learning and coordination, we introduced episodic memory, storing subgoal-level experiences for retrieval, and working memory, sharing environment-specific observations across nodes. Experiments on WAH-NL and ALFRED under partial observability show that ReAcTree consistently outperforms baselines across diverse LLMs, demonstrating improved reliability, generalization to unseen environments, and even enabling smaller models to rival larger ones.
@@
{acks}
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2024-00336738, Development of Complex Task Planning Technologies for Autonomous Agents, 40% and RS-2022-II220951, Development of Uncertainty-Aware Agents Learning by Asking Questions, 40%), and the National Research Council of Science & Technology (NST) grant by the Korea government (MSIT) (No. GTL25041-000, 20%).